OpenStudy (anonymous):

I need some help figuring out where I'm going wrong in calculation the questions is....The isotope^17f has 65s half life. If you start with a 15g sample of ^17f, after 32.5 s what is the remaining amount?

OpenStudy (anonymous):

So I know that N(t)/No=e (-0.693)32.5/65 but my answer is -0.3465 The answer is given as 0.707158819 please tell me what I've done wrong

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Hi yes can u see what I did wrong

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\huge N(t) = N _{0}*(\frac{ 1 }{2 })^{\frac{ t }{ 65 }}\]

OpenStudy (anonymous):

Ohh thankyou

OpenStudy (anonymous):

if you want to do it in terms of 'e' : \[N(t) = Ne ^{rt}\] \[\frac{ 1 }{ 2} = e ^{r65}\] \[\ln \frac{ 1 }{ 2} = r65\] \[\frac{ 1 }{ 65 } \ln \frac{ 1 }{ 2} = r\]

OpenStudy (anonymous):

r= -.010664

OpenStudy (anonymous):

either way you get 15g*.70710678...

OpenStudy (anonymous):

I don't know what I'm doing wrong but I can not get the 0.707158819

OpenStudy (anonymous):

0.707158819*No

OpenStudy (anonymous):

answer isn't .7071588 grams...

OpenStudy (anonymous):

The rest of the sum I get but I can't get the initial 0.707158819 to finish the rest of the sum

OpenStudy (anonymous):

this is less than a half-life, so the sample is even going to reach half it's original mass...

OpenStudy (anonymous):

sum?

OpenStudy (anonymous):

just plug 32.5 in for t

OpenStudy (anonymous):

32.5/65 = 1/2 (1/2)^(1/2) =....?

OpenStudy (anonymous):

Bugger me that was easier than I thought thank you Now it's just. 0.707158819*15*10^-3kg, awesome